40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 31 17:30:01 2022, Yehonathan, Update, BHD, Green shutters fixed 
    Reply  Wed Jun 1 11:56:28 2022, yuta, Update, ALS, Shutter driver for GRY replaced Screenshot_2022-06-01_12-01-53.png
Message ID: 16883     Entry time: Tue May 31 17:30:01 2022     Reply to this: 16884
Author: Yehonathan 
Type: Update 
Category: BHD 
Subject: Green shutters fixed 

[Paco, Yehonathan]

We fixed the slow control over the green beam shutters.

At the Y arm the wrong BNC was connected to the shutter driver. We connected the correct BNC to the driver and switched the remote mode. The green Y shutter now works but in reverese, meaning that sending 1 to C1:AUX-GREEN_Y_Shutter closes the shutter and vice versa. This needs to be fixed.

At the X end the problem was a bit more complicated. Previously, the shutter was controlled by c1auxey. We figured that c1auxex has a lot of spare bio channels. We found an Acromag BNC front panel (with wires already soldered to the BNCs) lying around in the lab and installed it on the c1auxex Acromag chassie. We then connected the topmost BNC to channel 0 on XT1111A in the chassie. The BNC was connected to the green shutter driver on the X end.

EPIC channel was added to the c1auxex db file while it was commented out on the psl shutters db file. Modbus was restarted on c1auxex and c1psl. c1psl had to be burt restored to regain MC lock. Now the green X shutter works properly.

ELOG V3.1.3-